OpenStudy (anonymous):

From home, Mary’s work is two thirds along the way to training. Training is 2.5km from work. Mary normally goes to work, then training and then home again. However, today she forgot her shoes. How far will Mary travel in total today if she has to go home before training to get her shoes?

OpenStudy (kainui):

Draw a picture and label how far the distance is, then use your pencil to trace the path out, remembering that you are following her same mistakes. then add up the total distance.

OpenStudy (anonymous):

let x = home to training 2x/3 = home to work x/3 = 2.5 km work to training so 2x/3 = 2 (2.5) = 5 today's travel: home to work 5 + work to home 5 + home to training (5 + 2.5) and back (5 + 2.5) = 25 km is it right??

OpenStudy (anonymous):

???